A bike successfully made a turn around a circular path whose radius is 13 m. At what angle is the road banked if it’s velocity was 12.0 m/s?


1
Expert's answer
2021-11-29T11:38:02-0500

Let's apply the Newton's Second Law of Motion in projections on "x"- and "y"-axis:


"Nsin\theta=\dfrac{mv^2}{r},""Ncos\theta=mg."

Dividing the first equation by the second one, we get:


"tan\theta=\dfrac{v^2}{gr},""\theta=tan^{-1}(\dfrac{v^2}{gr}),""\theta=tan^{-1}(\dfrac{(12.0\ \dfrac{m}{s})^2}{9.8\ \dfrac{m}{s^2}\times13\ m})=48.5^{\circ}."